OpenStudy (anonymous):

Solve for x. 2(–2 + 2x) = 4(x – 3)

OpenStudy (aum):

2(–2 + 2x) = 4(x – 3) distribute: -4 + 4x = 4x - 12 subtract 4x and you get -4 is equal to -12 which is never true. That implies there is NO solution to this equation. There is no value of x that will satisfy this equation.